# Joint Demonstration and Preference Learning Improves Policy Alignment with Human Feedback

### v2

@@ 6.1 MuJoCo Tasks
Results. We show that AIHF is able to integrate (insufficient amount of) demonstration data and (not-so-high-quality) preference data to generate high-quality policy, and it significantly outperforms the RLHF. In Fig. 3, we observe that due to the limited number of demonstration data Ross and Bagnell (2010); Zeng et al. (2022b), BC fails to provide a high-performing initialization for finetuning for RLHF. Moreover, since the preference data quality is only of low-to-medium quality, the RL step based on the learned reward model fails to significantly boost the fine-tuning performance. In contrast, clearly the proposed AIHF can effectively integrate preferences and demonstrations, leading to a more robust reward function and consequently, a high-quality policy.
6.2 LLM Task
In the experiment utilizing the HH dataset, we aim to generate helpful continuations through aligning large language models with human preference. We compare our proposed algorithm AIHF which updates reward and policy iteratively, with the standard RLHF benchmark. Additionally, $L_{2}(\cdot)$ is chosen to be $L_{2}^{\rm BTL}$ defined (11), and $L_{1}$, $L_{3}$ are chosen similarly as in MuJoCo experiment. More settings are listed in the Appendix A.2. For evaluation, we use the existing reward model PKU-Alignment/beaver-7b-v3.0-reward [^2] https://huggingface.co/PKU-Alignment/beaver-7b-v3.0-reward for the helpful continuation generation; use the existing sentiment classifier lvwerra/distilbert-imdb [^3] https://huggingface.co/lvwerra/distilbert-imdb as the ground truth reward model for the movie review generation task.
Data. We divide the training dataset of HH dataset into three different subsets: demonstration dataset, preference dataset, and RL training dataset, We employed the reward model PKU-Alignment/beaver-7b-v3.0-reward Dai et al. (2023); Ji et al. (2024) to filter the datasets. Specifically, We selected the 10k responses with the highest reward score as our demonstration dataset and utilized an additional 10,000 pairwise comparisons as preference data. The remaining prompts were then used to run the RL algorithm.
Setting. For the HH dataset, We first fine-tune the language models (Pythia-160M / Pythia-1B) Biderman et al. (2023) through supervised fine-tuning over all chosen responses from the HH dataset for 1 epoch, we call it full-SFT model and use it as our base model. Moreover, we also SFT the language model using the selected top 10k chosen responses and name it as demonstration-SFT model. For the reward model, we use EleutherAI/pythia-1.4B  [^4] https://huggingface.co/EleutherAI/pythia-1.4b for both 160m and 1B policy as its base model.
Figure 4: Helpfulness-controlled Generation. Left: Results on 160M models, Right: Results on 1B models. We record the average scores of AIHF and RLHF on the Anthropic-HH test dataset, reporting the results across three different trials.
Results. We observe that the proposed approach AIHF performs effectively when initiated from both the demonstration-SFT model and the full-SFT model. As shown in Fig. 4, utilizing the same data, the AIHF  algorithms can eventually outperform RLHF irrespective of the initial model. Furthermore, according to the numerical results as shown in Fig. 5, compared with the RLHF benchmark, we see that the proposed AIHF algorithm has smaller deviation from the base model. This benefit of the AIHF approach is due to the fact that we incorporate the maximum likelihood IRL objective for both reward learning and policy learning. In this case, both reward model and policy model will be trained to align with the demonstrations, which are also used in the training process of the SFT stage. Fig. 5 shows AIHF  effectively maintains the policy close to the initial model. This advantage can be partially attributed to Eq. (8a), as the inclusion of the demonstration regularization term encourages the alignment with the demonstration distribution, thus bringing better performance without significant deviation from the initial model. The win rate (ratio of samples where the reward of our model’s generation is higher than the model compared) provided by Fig 5 further shows the superiority of the proposed approach.
Figure 5: Helpfulness-controlled Generation on 1B models. Left: KL divergence to the Demonstration-SFT policy, Right: win rate of AIHF and RLHF over the Demonstration-SFT model.
7 Conclusion
[…]

### v3

@@ A.1.1 MuJoCo Tasks
Results. We show that AIHF is able to integrate (insufficient amount of) demonstration data and (not-so-high-quality) preference data to generate high-quality policy, and it significantly outperforms the RLHF and IPL. Note that IPL (Hejna & Sadigh, 2024) is actually the DPO-type algorithm applied to multi-horizon MDP, therefore it is an ideal baseline for the MuJoCo setting (since the underlying problem is a multi-horizon MDP). In Figure 5, We can see that AIHF outperforms both RLHF and IPL algorithms. We also observe that due to the limited number of demonstration data, even BC could provide a high-performing initialization, RLHF still fails to improve policy quality in the following RL stage Ross & Bagnell (2010); Zeng et al. (2022b). Moreover, since the preference data quality is only of low-to-medium quality, the RL step based on the learned reward model fails to significantly boost the fine-tuning performance. In contrast, clearly the proposed AIHF can effectively integrate preferences and demonstrations, leading to a more robust reward function and consequently, a high-quality policy.
The quality of the preference dataset and demonstration dataset are listed in Tab. 1.
Task Dataset Non-prefer Data Prefer Data Demonstration Data
Hopper-v2 $2345.20\pm 329.93$ $3024.63\pm 40.52$ $3559.61\pm 73.12$
Walker2d-v2 $3952.60\pm 444.45$ $5091.71\pm 291.73$ $5453.41\pm 71.07$
Table 1: The quality of preference and demonstration.
A.1.2 Sentiment-Controlled Generation
Dataset Generation: In the IMDb sentiment completion task, we generate the demonstrations and preference datasets using the following procedure. Initially, we train a Language Model by employing the ground-truth reward function distilbert-imdb and the Proximal Policy Optimization (PPO) algorithm on 30% of the training dataset for IMDb. Throughout the training process, we save the policy checkpoint every 500 PPO steps. Subsequently, we select an additional 40% of the training dataset and generate a response for each prompt for each checkpoint. According to the evaluation score of each generation, we categorize the data collected into different classes: low-, medium-, and high-quality datasets, then we combine low-quality and medium-quality as preference datasets, and use high-quality as demonstration datasets.
Training: After acquiring the preference and demonstration datasets, we train the proposed algorithm AIHF and baselines on the remaining 30% of prompts from the training dataset. We evaluate the performance of each algorithm using the test datasets for IMDb and HH, along with their corresponding ground truth reward functions. For the GPU resources, we use 8$\times$ A100 40G for all the experiments.
Results: Policy Quality. We find that the proposed approach works well when either preference or demonstration data, or both, are limited. From the 7, we see that by using the same amount of data (10k preference, 10k demonstration), AIHF-based algorithms achieve faster convergence than their RLHF and DPO counterparts.
A.1.3 The result of 1B and 2.8B experiments
Additional Results. We show here the distribution of the reward values of the continuation generated by the reward models trained by AIHF and RLHF. In Fig. 8, we can observe the distribution of AIHF and RLHF have overlaps in low-quality continuation, however, AIHF can generate more high-quality continuations compared to RLHF, which shows that joint optimization can more effectively align the policy model with the demonstration distribution.
We record the performance of the two variants, namely Direct AIHF (10) and Self-Play-AIHF (3.3), also RLHF, DPO, IPO and SPIN in Tab. 2. Our proposed AIHF still outperforms all these methods in this experiment setting.
Dataset DPO SPIN IPO RLHF Direct AIHF Self-Play-AIHF AIHF
10k Demonstrations, 5k preferences $0.463\pm 0.093$ $0.625\pm 0.048$ $0.616\pm 0.076$ $0.710\pm 0.085$ $0.752\pm 0.036$ $0.640\pm 0.102$ $1.167\pm 0.157$
10k Demonstrations, 10k preferences $0.474\pm 0.052$ $0.625\pm 0.048$ $0.650\pm 0.017$ $0.896\pm 0.095$ $0.798\pm 0.026$ $0.674\pm 0.096$ $1.190\pm 0.069$
Table 2: Evaluated reward scores of fine-tuned 1B Pythia models in HH dataset, evaluated by PKU-Alignment/beaver-7b-v3.0-reward.
A.1.4 The result of 7B experiments
We utilize DeepSpeed ZeRO-3 (Rajbhandari et al., 2020) to reduce the memory cost, and we use VLLM (Kwon et al., 2023) for accelerate data generation/inference. We use eight NVIDIA A100-40G to do the training with a per-device batch size of 1 for 7B model. We train all models with bfloat16 precision. We set the learning rate to be 5e-7 for the 7B model with the cosine learning rate scheduler, and the max sequence length is set to be 512.
Our device cannot conduct the implementation of the standard AIHF (5) using PPO. Therefore, we employed an online DPO approach (Xiong et al., 2024) for the lower-level policy optimization: 1) at each policy optimization iteration, we first generate multiple continuations from the current policy model over the given prompt dataset, 2) then we utilize our estimated explicit reward models to score each continuation, 3) over each prompt, choose the generated continuation with the highest reward score and the one with lowest reward score as a generated preference pair to run DPO algorithm to fine-tune the policy model.
For Self-Play AIHF (3.3) implementation, we adopt the same strategy as Chen et al. (2024), at each epoch, we generate samples with picked 50k data and generate continuation $\tilde{a}\sim\pi(\cdot|s)$ using the current model $\pi$, then optimize (3.3) with the sampled $\tilde{a}$.
To evaluate the effectiveness of our approach, we assessed our fine-tuned models on the widely used HuggingFace Open LLM Leaderboard (Beeching et al., 2023) in Table 3. We also list the metric and number of shots used for LLM evaluation on each dataset in Table 4. Moreover, in Tab. 5, we also utilize the Reward-Bench (Lambert et al., 2024) to evaluate the explicit reward models (estimated by standard preference learning in (3) and AIHF) and the implicit reward model (extracted from DPO). The implicit reward is calculated by the policy model similar to what the DPO paper Rafailov et al. (2023) proposed, i.e. $r(a|s):=\log(\frac{\pi(a|s)}{\pi^{0}(a|s)})$, the first explicit reward is estimated by BTL reward model (3), and second explicit reward is estimated by AIHF. The evaluation results show that AIHF can outperform the benchmark methods in terms of performance for both policy models and reward models. We also conduct an ablation study with different choices of $w_{1}$ in (10), as shown in Tab. 6, the improvement of joint learning methods over baseline is robust.
Dataset Arc Challenge TruthfulQA MC2 Winogrande GSM8K HellaSwag MMLU
Metric acc acc acc strict-match acc_norm acc
Num. of Shots 25 0 5 5 10 5
Table 3: A summarization of the benchmarks we use in this work. We list the metric and number of shots used for LLM evaluation on each dataset.
Tasks Arc Challenge TruthfulQA MC2 Winogrande GSM8k HellaSwag MMLU Avg
mistral-7b-sft-beta 54.69% 42.96% 77.27% 39.88% 82.23% 59.72% 59.46%
zephyr-7b-beta 59.64% 55.18% 77.82% 33.51% 84.19% 59.76% 61.68%
SPIN 58.45% 43.66% 78.30% 39.50% 83.59% 58.60% 60.35%
DPO 62.80% 53.17% 79.40% 39.20% 85.13% 59.41% 63.19%
IPO 58.02% 48.29% 79.24% 42.91% 83.93% 60.07% 62.08%
Direct AIHF 61.17% 60.03% 79.00% 39.80% 85.71% 60.02% 64.29%
Self-play AIHF 61.77% 58.29% 78.53% 44.20% 85.53% 58.66% 64.50%
AIHF 63.90% 58.38% 79.24% 40.56% 86.23% 60.18% 64.75%
Table 4: Test performance of Direct AIHF and Self-Play AIHF based on mistral-7b-sft-beta across HuggingFace Open LLM Leaderboard datasets.
Reward Model Chat Chat Hard Safety Reasoning Average
DPO Reward Model 37.43% 55.92% 64.14% 47.33% 51.21%
BTL Reward Model 95.11% 56.58% 63.69% 69.22% 71.15%
AIHF Reward Model 94.41% 55.37% 63.98% 76.75% 72.63%
Table 5: Evaluation of Reward Models in Reward-Bench.
Tasks Arc Challenge TruthfulQA MC2 Winogrande GSM8k HellaSwag MMLU Avg
Direct AIHF($w_{1}=0.01$) 61.86% 57.55% 79.08% 36.61% 85.58% 60.09% 63.46%
Direct AIHF($w_{1}=0.001$) 63.25% 58.73% 79.16% 36.84% 85.59% 59.26% 63.80%
Direct AIHF($w_{1}=0.0001$) 61.17% 60.03% 79.00% 39.80% 85.71% 60.02% 64.28%
Table 6: Test performance of Direct AIHF over different choices of the balancing coefficient across HuggingFace Open LLM Leaderboard datasets.
A.2 Why AIHF Can Outperform Two-Stage Alignment Approaches
In this section, we provide detailed analysis for Sec. 3.4.
A.2.1 SFT and RLHF Policy
We revisit the RLHF pipeline in the context of a simple softmax choice model where $\tau_{i}\in A$ is selected with probability
$\pi^{*}_{i}(R)=\frac{\exp(R_{i}/\beta)}{\sum_{j=1}^{N}\exp(R_{j}/\beta)}.$ (21)
For simplicity, we have defined $R_{i}:=R(\tau_{i})$. Using this policy model, below let us analyze different policies according to different ways that the rewards are learned.
Supervised Fine-Tuning (SFT): Given a demonstration dataset $\mathcal{D}$ the goal is the find the policy $\pi_{\rm SFT}$ that maximizes likelihood, i.e.:
$\pi_{\rm SFT}$ $:=\arg\max_{\pi}\sum_{\ell=1}^{N}\mathbb{E}_{\tau_{\ell}\sim\mathcal{D}}\left[\log\pi^{*}_{\ell}(R)\right],\quad\mbox{\rm s.t.}\;\sum_{\ell=1}^{N}\pi^{*}_{\ell}(R)=1.$ (22)
Next let us identify the optimal reward and its corresponding policy in this case. First, let us find the reward function. Write (22) in terms of reward optimization, we have:
$R^{*}=\arg\max_{R}\sum_{\ell=1}^{N}\mathbb{E}_{\tau_{\ell}\sim D}\log\frac{\exp(R_{\ell}/\beta)}{\sum_{j=1}^{N}\exp(R_{j}/\beta)}=:L_{1}(R)$ (23)
The partial derivative of the objective function w.r.t. a component $R_{i},i\in[1,\cdots,N]$ is given by:
$\frac{\partial L_{1}(R)}{\partial R_{i}}$ $=\sum_{\ell:\ell\neq i}\mathbb{E}_{\tau_{\ell}\sim D}\left[-\frac{1}{\beta}\frac{\exp(R_{i}/\beta)}{\sum_{j=1}^{N}\exp(R_{j}/\beta)}\right]+\mathbb{E}_{\tau_{i}\sim D}\left[\frac{1}{\beta}-\frac{1}{\beta}\frac{\exp(R_{i}/\beta)}{\sum_{j=1}^{N}\exp(R_{j}/\beta)}\right]$
$=\frac{1}{\beta}\left(\frac{\#\{\tau_{i}\in\mathcal{D}\}}{|P|}-\sum_{\ell}\mathbb{E}_{\tau_{\ell}\in D}\left[\frac{\exp(R_{i}/\beta)}{\sum_{j=1}^{N}\exp(R_{j}/\beta)}\right]\right),\quad i\in[1:N].$ (24)
Setting the above partial gradient to zero, we obtain
#$\frac{\#\{\tau_{i}\in\mathcal{D}\}}{|\mathcal{D}|}=\frac{\exp(\widehat{R}_{\mathcal{D}}(\tau_{i})/\beta)}{\sum_{j=1}^{N}\exp(\widehat{R}_{\mathcal{D}}(\tau_{j})/\beta)}~{}~{}~{}~{}i\in[1:N],$ (25)
where we used $\widehat{R}_{D}(\tau_{i})$’s to denote the optimal reward estimated from the set of SFT data $\mathcal{D}$. Then according to the definition of the policy given earlier, we obtain
$\pi^{*}_{i}(\widehat{R}_{\mathcal{D}})=\frac{\#\{\tau_{i}\in\mathcal{D}\}}{|\mathcal{D}|},\;\forall~{}i\in[1,\cdots N].$
Note that the optimal reward learned from the SFT data $\widehat{R}_{D}$ should satisfy the system (25). It turns out that there is a unique solution to the following system of equations:
#$\frac{\#\{\tau_{i}\in\mathcal{D}\}}{|\mathcal{D}|}=\frac{\exp(\widehat{R}_{\mathcal{D}}(\tau_{i})/\beta)}{\sum_{j=1}^{N}\exp(\widehat{R}_{\mathcal{D}}(\tau_{j})/\beta)}~{}~{}~{}~{}i\in\{2,\dots,N\}$ (26)
with $\widehat{R}_{\mathcal{D}}(\tau_{1})=\bar{R}_{1}$ a fixed reference value. A simple argument is provided below. The system (26) can be re-written as:
$\log\frac{\#\{\tau_{i}\in\mathcal{D}\}}{|\mathcal{D}|}-\log\frac{\#\{\tau_{1}\in\mathcal{D}\}}{|\mathcal{D}|}$ $=\frac{1}{\beta}\big{(}\widehat{R}_{\mathcal{D}}(\tau_{i})-\bar{R}_{1}\big{)}~{}~{}~{}~{}~{}i\in\{2,\dots,N\}$
Hence, rewards are uniquely determined by demonstration data with a fixed reference value $\widehat{R}_{\mathcal{D}}(\tau_{1})=\bar{R}_{1}$.
It is important to note that, despite the fact that in the SFT we may not directly learn a reward function, the above analysis says that we are implicitly learning a reward $\widehat{R}_{D}$, assuming that the policy is parameterized as (21). This implicit reward will be used later to analyze the RLHF policy. Indeed, if we solve the following SFT policy optimization problem directly
$\max\sum_{\ell}\mathbb{E}_{\tau_{\ell}\in\mathcal{D}}\log(\pi_{\ell}),\quad{\rm s.t.}\quad\sum_{i\in\mathcal{D}}\pi_{i}=1,$ (27)
one can easily obtain that the optimal policy is given by $\pi^{*}_{i}=\frac{\#\{\tau_{i}\in\mathcal{D}\}}{|\mathcal{D}|},\;\forall~{}i\in[1,\cdots N]$.
Reward Modeling with Preference Data: Now let us analyze the case where the reward is only learned from a preference dataset. With preference data $\mathcal{P}:=\{(\tau_{i}\succ\tau_{j})\}$, the BTL model is:
$P(\tau_{i}\succ\tau_{j})$ $=\sigma\big{(}\frac{1}{\beta}\big{(}R(\tau_{i})-R(\tau_{j})\big{)}=\frac{\pi^{*}_{i}(R)}{\pi^{*}_{i}(R)+\pi^{*}_{j}(R)}.$
The reward estimation problem is defined as:
$\widehat{R}_{\mathcal{P}}=\arg\max_{R}L_{2}(R)$ $:=\sum_{\ell,j:\ell\neq j}\mathbb{E}_{(\tau_{\ell}\succ\tau_{j})\sim\mathcal{P}}\Big{[}\log\frac{\pi^{*}_{\ell}(R)}{\pi^{*}_{\ell}(R)+\pi^{*}_{j}(R)}\Big{]}$ (28)
$=\sum_{\ell,j:\ell\neq j}\mathbb{E}_{(\tau_{\ell}\succ\tau_{j})\sim\mathcal{P}}\left[\log\left(\frac{\exp(R_{\ell}/\beta)}{\exp(R_{\ell}/\beta)+\exp(R_{j}/\beta)}\right)\right]$ (29)
Take the gradient w.r.t. $R_{i}$, we obtain
$\frac{\partial\ell_{RM}({R})}{\partial R_{i}}$ $=\sum_{j:j\neq i}\mathbb{E}_{(\tau_{i}\succ\tau_{j})}\left(\frac{1}{\beta}-\frac{1}{\beta}\left(\frac{\exp(R_{\ell}/\beta)}{\exp(R_{\ell}/\beta)+\exp(R_{j}/\beta)}\right)\right)$
$\quad+\sum_{j:j\neq i}\mathbb{E}_{(\tau_{j}\succ\tau_{i})}\left[-\frac{1}{\beta}\left(\frac{\exp(R_{\ell}/\beta)}{\exp(R_{\ell}/\beta)+\exp(R_{j}/\beta)}\right)\right]$
$=\frac{1}{\beta}\sum_{j:j\neq i}\frac{|\mathcal{P}_{i\succ j}|}{|\mathcal{P}|}\left(1-\left(\frac{\exp(R_{\ell}/\beta)}{\exp(R_{\ell}/\beta)+\exp(R_{j}/\beta)}\right)\right)$
$\quad+\frac{1}{\beta}\sum_{j:j\neq i}\frac{|\mathcal{P}_{j\succ i}|}{|\mathcal{P}|}\left(-\left(\frac{\exp(R_{\ell}/\beta)}{\exp(R_{\ell}/\beta)+\exp(R_{j}/\beta)}\right)\right)$
$=\frac{1}{\beta}\left(\sum_{j:j\neq i}\frac{|\mathcal{P}_{i\succ j}|}{|\mathcal{P}|}-\frac{|\mathcal{P}_{i,j}|}{|\mathcal{P}|}\frac{\exp(R_{\ell}/\beta)}{\exp(R_{\ell}/\beta)+\exp(R_{j}/\beta)}\right)$
$=\frac{1}{\beta}\sum_{j:j\neq i}\Big{(}\frac{|\mathcal{P}_{i\succ j}|}{|\mathcal{P}_{i,j}|}-\frac{\pi^{*}_{i}(R)}{\pi^{*}_{i}(R)+\pi^{*}_{j}(R)}\Big{)}\frac{|\mathcal{P}_{i,j}|}{|\mathcal{P}|}$ (30)
where in the above derivation we have defined $|\mathcal{P}_{i\succ j}|:=\#\{\tau_{i}\succ\tau_{j}~{}\mbox{in}~{}\mathcal{P}\}$ and $|\mathcal{P}_{i,j}|:=|\mathcal{P}_{i\succ j}|+|\mathcal{P}_{j\succ i}|$. Setting the gradient to zero, we obtain
$\sum_{j:j\neq i}\Big{(}\frac{|\mathcal{P}_{i\succ j}|}{|\mathcal{P}_{i,j}|}-\frac{\pi^{*}_{i}(\widehat{R}_{P})}{\pi^{*}_{i}(\widehat{R}_{P})+\pi^{*}_{j}(\widehat{R}_{P})}\Big{)}\frac{|\mathcal{P}_{i,j}|}{|\mathcal{P}|}=0.$ (31)
We shall denote by $\widehat{R}_{\mathcal{P}}$ as a solution to the above system of equations. The first-order condition (31) can be written in implicit form as:
$\pi^{*}_{i}(\widehat{R}_{P})=\frac{\sum_{j:j\neq i}|\mathcal{P}_{i\succ j}|}{\sum_{j:j\neq i}\frac{|\mathcal{P}_{ij}|}{\pi_{i}^{*}(\widehat{R}_{P})+\pi_{j}^{*}(\widehat{R}_{P})}}=\frac{\sum_{j:j\neq i}|\mathcal{P}_{i\succ j}|}{\sum_{j:j\neq i}\frac{|\mathcal{P}_{ij}|}{1-\sum_{\ell\in A\backslash\{i,j\}}\pi^{*}_{\ell}(\widehat{R}_{P})}}:=\frac{\sum_{j:j\neq i}|\mathcal{P}_{i\succ j}|}{\sum_{j:j\neq i}{|\mathcal{P}_{ij}|}\rho_{-(i,j)}(\pi^{*}(\widehat{R}_{P}))}$ (32)
where we have defined $\rho_{-(i,j)}(\pi):=\Big{(}1-\sum_{\ell\in A\backslash\{i,j\}}\pi_{\ell}\Big{)}^{-1}$ as the expected number of times an action other than $\tau_{i}$ or $\tau_{j}$ is selected when sampling actions from $\pi$ infinitely many times.
RLHF Policy. Based on the above two analysis, we are ready to analyze the RLHF policy. The RLHF policy is defined as follows:
where $\widehat{R}_{\mathcal{P}}$ is the estimator obtained from preference data, $\pi^{\rm SFT}$ is the SFT model trained with demonstration dataset $\mathcal{D}$, and $\Delta^{N}$ is the probability simplex.
It can be easily shown that the solution $\pi^{\rm RLHF}$ is of the form:
$\pi^{\rm RLHF}(\tau_{i})$ $=\frac{\pi^{\rm SFT}(\tau_{i})\exp\left(\frac{1}{\beta}\widehat{R}_{\mathcal{P}}(\tau_{i})\right)}{\sum_{j=1}^{N}\pi^{\rm SFT}(\tau_{j})\exp\left(\frac{1}{\beta}\widehat{R}_{\mathcal{P}}(\tau_{j})\right)}$
[…]
$=\pi^{*}_{i}\Big{(}\widehat{R}_{\mathcal{D}}+\widehat{R}_{\mathcal{P}}\Big{)}.$ (33)
A.2.2 The AIHF Policy
The AIHF Policy. The proposed AIHF estimation problem is
$\widehat{R}^{\rm AIHF}=\arg\max_{R}\ell_{\mathcal{D}+\mathcal{P}}(R):=|\mathcal{D}|L_{1}(R)+|\mathcal{P}|L_{2}(R)$ (34)
where $L_{1}(R):=\mathbb{E}_{\tau_{i}\sim\mathcal{D}}[\log\pi^{*}_{i}(R)]$ and $L_{2}(R):=\mathbb{E}_{(\tau_{j}\prec\tau_{i})\sim\mathcal{P}}[\log\frac{\pi^{*}_{i}(R)}{\pi^{*}_{i}(R)+\pi^{*}_{j}(R)}]$, and $\pi^{*}_{i}(R)$ is given by (21).
Similarly as before, taking gradient w.r.t. $R_{i}$, and leverage (A.2.1) and (A.2.1), we obtain
$\frac{\partial\ell_{\mathcal{D}+\mathcal{P}}(R)}{\partial R_{i}}$ =#$=\frac{\#\{\tau_{i}~{}\mbox{in}~{}\mathcal{D}\}}{|\mathcal{D}|}|\mathcal{D}|-\pi^{*}_{i}(R)|\mathcal{D}|+\sum_{j:j\neq i}\left(|\mathcal{P}_{i\succ j}|-\sum_{j\neq i}\frac{\pi^{*}_{i}(R)}{\pi^{*}_{i}(R)+\pi^{*}_{j}(R)}|\mathcal{P}_{i,j}|\right)$ (35)
Setting the above condition to zero, we obtain that the AIHF reward should satisfy the following
#$\#\{\tau_{i}~{}\mbox{in}~{}\mathcal{D}\}+\sum_{j:j\neq i}|\mathcal{P}_{i\succ j}|$ $=\pi^{*}_{i}(\widehat{R}^{\rm AIHF})\Big{(}|\mathcal{D}|+\sum_{j:j\neq i}\frac{|\mathcal{P}_{i,j}|}{\pi^{*}_{i}(\widehat{R}^{\rm AIHF})+\pi^{*}_{j}(\widehat{R}^{\rm AIHF})}\Big{)}$
$=\pi^{*}_{i}(\widehat{R}^{\rm AIHF})\Big{(}|\mathcal{D}|+\sum_{j:j\neq i}|\mathcal{P}_{i,j}|\rho_{-(i,j)}\big{(}\pi^{*}(\widehat{R}^{\rm AIHF})\big{)}\Big{)}$
Or equivalently,
$\pi^{*}_{i}(\widehat{R}^{\rm AIHF})$ =#$=\frac{\#\{\tau_{i}~{}\mbox{in}~{}\mathcal{D}\}+\sum_{j\neq i}|\mathcal{P}_{i\succ j}|}{|\mathcal{D}|+\sum_{j:j\neq i}|\mathcal{P}_{i,j}|\rho_{-(i,j)}\big{(}\pi^{*}(\widehat{R}^{\rm AIHF})\big{)}}$ (36)
The system (36) has a unique solution $\widehat{R}^{\rm AIHF}$ with a fixed reference value $\widehat{R}^{\rm AIHF}(\tau_{1})=\bar{R}_{1}$.
Discussion. Now let us compare the RLHF policy (33) and AIHF policy (36). First, observe that the RLHF policy is a function of the sum of two reward functions, one learned from the demonstration data and one learned from the preference data, while the AIHF policy takes a more complex form. To have a better understanding of the two policies, let us consider the extreme cases where the SFT and the preference datasets are unbalanced, where either $|\mathcal{D}|\gg|\mathcal{P}|$ or $|\mathcal{P}|\gg|\mathcal{D}|$.
  • Case $|\mathcal{D}|\gg|\mathcal{P}|$. In this case, $\pi^{*}_{i}(\widehat{R}^{\rm AIHF})\approx\frac{\#\{\tau_{i}~{}\mbox{in}~{}\mathcal{D}\}}{|\mathcal{D}|}=\pi^{*}_{i}(\widehat{R}_{\mathcal{D}})$, so the AIHF policy will ignore the preference dataset, while focusing on the SFT policy; however, for the RLHF policy, one needs to first separately performs the reward estimation, but in this case $\hat{R}_{P}$ can be very noisy due to lack of data; Then the RLHF policy will also be noisy since it will be influenced by the noisy reward function $\hat{R}_{P}$.
  • Case $|\mathcal{P}|\gg|\mathcal{D}|$. In this case, let’s further assume that $\sum_{j:j\neq i}|\mathcal{P}_{i\succ j}|\gg\#\{\tau_{i}~{}\mbox{in}~{}\mathcal{D}\},\;\forall~{}i$, that is, for each trajectory $\tau_{i}$, the number of times it appears in the preference data is much larger than its appearance in the demonstration data. In this case, it is easy to see that $\pi^{*}_{i}(\widehat{R}^{\rm AIHF})\approx\pi^{*}_{i}(\widehat{R}_{\mathcal{P}})$, while the RLHF policy will still be noisy as in the first case.
A.2.3 Numerical Examples
Example 1: With $\beta=1$ and only two actions $\tau_{1}$ and $\tau_{2}$. Since $\rho_{-(1,2)}(\pi)=\rho_{-(2,1)}(\pi)=1$, it follows from equations (26), (32), (36) that:
$\pi^{\rm AIHF}_{1}:=\pi^{*}_{1}(\widehat{R}^{\rm AIHF})$ =#$=\frac{\#\{\tau_{1}~{}\mbox{in}~{}\mathcal{D}\}+\#\{\tau_{2}\prec\tau_{1}~{}\mbox{in}~{}\mathcal{P}\}}{|\mathcal{D}|+|\mathcal{P}|}$
$=\frac{|\mathcal{D}|}{|\mathcal{D}|+|\mathcal{P}|}\pi^{*}_{1}(\widehat{R}_{\mathcal{D}})+\frac{|\mathcal{P}|}{|\mathcal{D}|+|\mathcal{P}|}\pi^{*}_{1}(\widehat{R}_{\mathcal{P}}).$
Slightly abusing notations, let $\pi^{*}_{1}:=\pi^{*}_{1}(R^{*})$ where $R^{*}$ is the ground-truth reward. It follows that $\mbox{Var}(\pi^{*}_{1}(\widehat{R}_{\mathcal{D}}))=\frac{\pi^{*}_{1}(1-\pi^{*}_{1})}{|\mathcal{D}|}$, $\mbox{Var}(\pi^{*}_{1}(\widehat{R}_{\mathcal{P}}))=\frac{\pi^{*}_{1}(1-\pi^{*}_{1})}{|\mathcal{P}|}$ and
$\mbox{Var}(\pi^{\rm AIHF}_{1})=\frac{\pi^{*}_{1}(1-\pi^{*}_{1})}{|\mathcal{D}|+|\mathcal{P}|}<\min\{\mbox{Var}(\pi^{*}_{1}(\widehat{R}_{\mathcal{D}})),\mbox{Var}(\pi^{*}_{1}(\widehat{R}_{\mathcal{P}}))\}$
Hence, the AIHF policy estimate has less variance than either the policy obtained from only demonstrations or preferences.
To further illustrate, suppose the ground truth is $R^{*}(\tau_{1})=R^{*}(\tau_{2})$ and in the dataset there are more demonstrations than preferences, i.e. $|\mathcal{D}|\gg|\mathcal{P}|$:
With the given data, $\pi^{\rm SFT}_{1}=\pi^{*}_{1}(\widehat{R}_{\mathcal{D}})=\frac{\#\{\tau_{1}~{}\mbox{in}~{}\mathcal{D}\}}{|\mathcal{D}|}=\frac{50}{100}$ and the solution to (32) yields
$\pi^{*}_{1}(\widehat{R}_{\mathcal{P}})=\frac{\exp\widehat{R}_{\mathcal{P}}(\tau_{1})}{\exp\widehat{R}_{\mathcal{P}}(\tau_{1})+\exp\widehat{R}_{\mathcal{P}}(\tau_{2})}=\frac{6}{10}.$
$=\frac{\exp\widehat{R}_{\mathcal{P}}(\tau_{1})}{\exp\widehat{R}_{\mathcal{P}}(\tau_{1})+\exp\widehat{R}_{\mathcal{P}}(\tau_{2})}=\frac{6}{10}.$
In this example, the RLHF policy estimator is the furthest away from ground-truth, because it does not correctly use the information provided by the demonstration data which in this case happens by chance to be correct $\pi^{\rm SFT}_{1}=\pi^{*}_{1}(\widehat{R}_{\mathcal{P}})=\frac{1}{2}$.
As a second example, again suppose the ground truth is $R^{*}(\tau_{1})=R^{*}(\tau_{2})$ and in the dataset there are more preferences than demonstrations, $|\mathcal{P}|\gg|\mathcal{D}|$:
{#$\{\#\tau_{1}~{}\mbox{in}~{}\mathcal{D}\}=6,\quad\{\#\tau_{2}~{}\mbox{in}~{}\mathcal{D}\}=4,\quad\{\#\tau_{1}\succ\tau_{2}~{}\mbox{in}~{}\mathcal{P}\}=\{\#\tau_{2}\succ\tau_{1}~{}\mbox{in}~{}\mathcal{P}\}=50.$ (37)
In this case, $\pi^{\rm SFT}_{1}=\pi^{*}_{1}(\widehat{R}_{\mathcal{D}})=\frac{6}{10}$ and the solution to (32) yields
$\pi^{*}_{1}(\widehat{R}_{\mathcal{P}})=\frac{\exp\widehat{R}_{\mathcal{P}}(\tau_{1})}{\exp\widehat{R}_{\mathcal{P}}(\tau_{1})+\exp\widehat{R}_{\mathcal{P}}(\tau_{2})}=\frac{50}{100}.$
It follows from (33) that:
$=\frac{\pi_{1}^{SFT}}{\pi_{1}^{SFT}+\pi_{2}^{SFT}}=\frac{6}{10}.$
Example 2: Let us use an illustrative example to show that RLHF method will result in significant data under-utilization when the demonstration coverage is limited. With $\beta=1$, assume that there are 50 actions, i.e. $A=\{1,2,\dots,50\}$ and each with a ground-truth reward defined by $R^{*}(\tau_{i})=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac{(\frac{i}{50}-\mu)^{2}}{2\sigma^{2}}}$, where $\mu=0.5$ and $\sigma=2$. Assume we can sample demonstration and preference from the ground truth reward distribution: demonstrations are sampled from the multinomial distribution, while preferences are sampled from the BTL model.
[…]